# On Function-Correcting Codes in the Lee Metric

### v1

@@ Abstract
Function-correcting codes are a coding framework designed to minimize redundancy while ensuring that specific functions or computations of encoded data can be reliably recovered, even in the presence of errors. The choice of metric is crucial in designing such codes, as it determines which computations must be protected and how errors are measured and corrected. Previous work by Liu and Liu [6] studied function-correcting codes over $\mathbb{Z}_{2^{l}},\ l\geq 2$ using the homogeneous metric, which coincides with the Lee metric over $\mathbb{Z}_{4}$. In this paper, we extend the study to codes over $\mathbb{Z}_{m},$ for any positive integer $m\geq 2$ under the Lee metric and aim to determine their optimal redundancy. To achieve this, we introduce irregular Lee distance codes and derive upper and lower bounds on the optimal redundancy by characterizing the shortest possible length of such codes. These general bounds are then simplified and applied to specific classes of functions, including Lee-local functions, Lee weight functions, and Lee weight distribution functions, leading to improved some bounds compared to those of Liu and Liu [6] over $\mathbb{Z}_{4}$ and generalize the other bounds over $\mathbb{Z}_{m}$ in the Lee metric.
Keywords: Function-correcting codes, error-correcting codes, Lee metric, redundancy.
@@ 1 Introduction
Lenz et al. [5] first introduced the framework of function-correcting codes (FCC) in the Hamming metric, establishing the foundational theory for correcting function evaluations rather than raw symbols. This framework was later extended to other metrics to capture different error models. Specifically, Xia et al. [15] generalized FCC to the symbol-pair metric, while Singh et al. [13] extended it further to the b-symbol metric. Building on these developments, in [14], the authors introduced locally b-symbol codes and investigated the redundancy of the corresponding function-correcting b-symbol codes. Several related advances complement these efforts: Ge et al. [2] studied optimal redundancy bounds for such codes, Ly et al. [7] proposed a conjecture concerning their asymptotic performance, and Premlal et al. [8] developed a function correcting framework for linear functions. In the context of locality, Rajput et al. [9] investigated locally function-correcting codes for the Hamming metric, while Sampath et al. [11] contributed further generalizations of FCC to specialized coding scenarios. Together, these works form a cohesive progression from classical FCC in the Hamming setting to rich variants supporting diverse metrics, locality, and function-specific error corrections.
@@
Function-correcting codes with respect to the homogeneous metric were proposed in [6], where the authors introduced function-correcting codes with homogeneous distance (FCCHDs) as a generalization of function-correcting codes under the Hamming metric. Since both Hamming weight and Lee weight over $\mathbb{Z}_{4}$ can be regarded as special cases of homogeneous weights, the results in [6] naturally extend to the Lee metric over $\mathbb{Z}_{4}$. Motivated by this fact, we investigate the construction and properties of function-correcting codes for the Lee metric over $\mathbb{Z}_{m}$, where $m\geq 2$ is an arbitrary positive integer. We derive upper and lower bounds on the optimal redundancy of function-correcting Lee metric codes for arbitrary functions by employing shortest-length irregular Lee distance codes in Section 3. In Section 4, we provide several bounds on the shortest-length irregular Lee distance codes analogous to the Plotkin and Gilbert-Vashamov bounds. As a result, we improve the bound derived in [6, Lemma 3.2] over $\mathbb{Z}_{4}$. Furthermore, we introduce the concept of locally Lee functions, and using general theoretical results, establish bounds on the optimal redundancy for this class of functions in Section 5. In particular, we demonstrate that the lower bound on optimal redundancy is achievable for locally binary Lee functions. Additionally, we investigate the locality properties of specific functions, the Lee weight function and the Lee weight distribution function, and determine bounds on their optimal redundancy in Section 6. The bounds on the optimal redundancy of function-correcting codes can be obtained as a direct consequence of our general results. In particular, by substituting $m=2$ and $m=4$ into Lemma 4.1, we get the specific cases corresponding to codes in the Hamming distance over $\mathbb{Z}_{2}$ in [5] and the homogeneous distance over $\mathbb{Z}_{4}$ in [6], respectively. This shows that the well-known bounds for these two settings follow naturally from our general framework, without requiring any additional arguments.
@@ 2 Preliminaries
In this section, we discuss basic definitions and results that will be used throughout the paper. We denote the residue ring modulo $m$ by $\mathbb{Z}_{m}$, where $m$ is a positive integer greater than or equal to $2$. We consider the elements of $\mathbb{Z}_{m}$ as $10,1,\dots,m-$ 1 unless otherwise specified. For a positive integer $k$, $\mathbb{Z}_{m}^{k}$ denotes the set of all $k$-tuples over $\mathbb{Z}_{m}$.
@@
$\begin{split}-\frac{m-1}{2},-\frac{m-3}{2},\dots,-1,0,1,\dots,\frac{m-1}{2}&\text{ if }m\text{ is odd}\\ -\frac{m-2}{2},\dots,-1,0,1,\dots,\frac{m-2}{2},\frac{m}{2}&\text{ if }m\text{is even},\end{split}$
@@ 3 Function-correcting Lee metric codes
Any classical error-correcting code with the minimum Lee distance $12t+$ 1, defines an $(f,t)$-FCLMC for any function $f:\mathbb{Z}_{m}^{k}\to Im(f)$. For an integer $n$, we denote $[n]=\{1,2,\dots,n\}$ and $[n]^{+}=\max\{n,0\}$. Define the distance matrix of a given function $f$ as follows.
@@
$[D_{f}^{L}(t,u_{1},u_{2},\dots,u_{M})]_{ij}=\begin{cases}[2t+1-d_{L}(u_{i},u_{j})]^{+}&\text{if }f(u_{i})\neq f(u_{j})\\ 0&\text{ otherwise}\end{cases}$
@@
Let $N_{L}(D)$ be the smallest integer $r$ such that there exists a $D$-code of length $r$. When $[D]_{ij}=d$ for all $i\neq j$, we denote $N_{L}(M,d)$. Now, we prove a fundamental relationship between the redundancy of optimal FCLMCs and the irregular distance codes.
@@
$[D_{f}^{L}(t,e_{1},e_{2},\dots,e_{\eta})]_{ij}=\begin{cases}[2t+1-d_{L}(e_{i},e_{j})]^{+}&\text{ if }i\neq j\\ 0&\text{otherwise}.\end{cases}$
@@ 4 Bounds on the shortest length of irregular-distance codes
$N_{L}(D)\geq\begin{cases}\frac{4}{M^{2}\left\lfloor\frac{m}{2}\right\rfloor}\sum_{i,j;i<j}[D]_{ij},&\text{ if }M\text{ is even}\\ \frac{4}{(M^{2}-1)\left\lfloor\frac{m}{2}\right\rfloor}\sum_{i,j;i<j}[D]_{ij},&\text{ if }M\text{ is odd.}\end{cases}$
@@
Let $N_{L}(D)=r$. Let $C\subseteq\mathbb{Z}_{m}^{r}$ be a $D$-code of length $r$ and size $M$. Consider an $M\times r$ matrix $A$ whose rows are codewords $\{C_{1},C_{2},\dots,C_{M}\}=C$. First, we derive an upper bound on the sum of the Lee distances between all the codewords.
$\sum_{x,y\in C}d_{L}(x,y)=\sum_{x,y\in C}\sum_{j=1}^{r}d_{L}(x_{j},y_{j})=\sum_{j=1}^{r}\sum_{x,y\in C}d_{L}(x_{j},y_{j}).$
Now, we calculate $\sum_{x,y\in C}d_{L}(x_{j},y_{j}).$ Let $A_{j}$ be the $j^{\mathrm{th}}$-column of $A$. Let $n_{\alpha}$ denote the number of $\alpha\in\mathbb{Z}_{m}$ in $A_{j}$. If $M$ is even, then
$\sum_{x,y\in C}d_{L}(x_{j},y_{j})=\sum_{\alpha,\beta\in\mathbb{Z}_{m}}n_{\alpha}n_{\beta}d_{L}(\alpha,\beta)\leq 2\cdot\frac{M^{2}}{4}\left\lfloor\frac{m}{2}\right\rfloor.$
Maximum attained when $n_{\alpha}=M/2=n_{\beta}$ such that $d_{L}(\alpha,\beta)=\lfloor\frac{m}{2}\rfloor$ for some $\alpha,\beta\in\mathbb{Z}_{m}$ and $n_{\gamma}=0$ for other $\gamma\in\mathbb{Z}_{m}$. Therefore, we have
$\sum_{x,y\in C}d_{L}(x,y)\leq r\frac{M^{2}}{2}\left\lfloor\frac{m}{2}\right\rfloor\ \text{i.e.}\sum_{1\leq i,j\leq N}d_{L}(C_{i},C_{j})\leq r\frac{M^{2}}{2}\left\lfloor\frac{m}{2}\right\rfloor$
@@
$r\geq\frac{2}{M^{2}\left\lfloor\frac{m}{2}\right\rfloor}\sum_{i,j}[D]_{ij}=\frac{4}{M^{2}\left\lfloor\frac{m}{2}\right\rfloor}\sum_{i,j;i<j}[D]_{ij}.$
When $M$ is odd, then
$\sum_{x,y\in C}d_{L}(x_{j},y_{j})=\sum_{\alpha,\beta\in\mathbb{Z}_{m}}n_{\alpha}n_{\beta}d_{L}(\alpha,\beta)\leq 2\cdot\frac{(M-1)(M+1)}{4}\left\lfloor\frac{m}{2}\right\rfloor.$
The rest proof follows similarly. ∎
In the context of regular-distance codes possessing a minimum distance of $d$, by Lemma 4.1, it follows that
$N(M,d)\geq\frac{2d(M-1)}{M\left\lfloor\frac{m}{2}\right\rfloor}.$
@@
Let $m=4$. For any Lee distance matrix $D\in\mathbb{N}_{0}^{M\times M}$, we have
$N_{L}(D)\geq\begin{cases}\frac{2}{M^{2}}\sum_{i,j;i<j}[D]_{ij},&\text{ if }M\text{ is even}\\ \frac{2}{(M^{2}-1)}\sum_{i,j;i<j}[D]_{ij},&\text{ if }M\text{ is odd.}\end{cases}$
Remark 4.3.
As the homogeneous metric coincides with the Lee metric over $\mathbb{Z}_{4}$. In [6, Lemma 3.2], Liu et al. provided a lower bound on $N_{L}(D)$. The above Corollary 4.2 gives a better lower bound on $N_{L}(D)$ over $\mathbb{Z}_{4}$. Precisely, it improves the bound in [6] by a multiple of $2$ over $\mathbb{Z}_{4}$.
It is well known that the Lee metric coincides with the Hamming metric over $\mathbb{Z}_{2}$. Therefore, [5, Lemma 1] can be derived from Lemma 4.1 by putting $m=2$. We state it in the following corollary.
@@
Let $m=2$. For any Lee distance matrix $D\in\mathbb{N}_{0}^{M\times M}$, we have
$N_{L}(D)\geq\begin{cases}\frac{4}{M^{2}}\sum_{i,j;i<j}[D]_{ij},&\text{ if }M\text{ is even}\\ \frac{4}{(M^{2}-1)}\sum_{i,j;i<j}[D]_{ij},&\text{ if }M\text{ is odd.}\end{cases}$
@@
Let $D$ be a distance matrix with $[D]_{ij}=d$ for all $i\neq j$. Then $N_{L}(M,d)\leq N_{H}(M,d)$.
@@ 5 FCLMCs for the locally functions
Let $f:\mathbb{F}_{2}^{k}\to Im(f)$ be a locally $(\lambda,\rho)_{L}$-function. Then there exists a map $\tau_{f}^{L}:\mathbb{F}_{2}^{k}\to[\lambda]$ such that $Im(\tau_{f}^{L})\leq[\lambda]$ and for all $u,v\in\mathbb{F}_{2}^{k}$, $\tau_{f}^{L}(u)\neq\tau_{f}^{L}(v)$ if either $u\in B_{L}^{f}(v,\rho)$ or $v\in B_{L}^{f}(u,\rho)$.
@@
By Lemma 5.3, there exists a map $\tau_{f}^{L}:\mathbb{Z}_{m}^{k}\to[\lambda]$ such that for $u,v\in\mathbb{F}_{2}^{k}$, $\tau_{f}^{L}(u)\neq\tau_{f}^{L}(v)$ whenever $f(u)\in B_{L}^{f}(v,2t)$ or $f(u)\in B_{L}^{f}(u,2t)$. Let $C=\{C_{1},C_{2},\dots,C_{\lambda}\}$ be a code of length $N_{L}(\lambda,2t)$ with $\lambda$ codewords and minimum Lee distance $2t$. Define an encoding $\mathcal{E}:\mathbb{Z}_{m}^{k}\to\mathbb{Z}_{m}^{k+N_{L}(\lambda,2t)}$ given by $\mathcal{E}(u)=(u,C_{\tau_{f}^{L}(u)})$. We claim that this encoding defines an $(f,t)$-FCLMC. Let $u,v\in\mathbb{Z}_{m}^{k}$ with $f(u)\neq f(v)$. Then
@@
Case 2: If $d_{L}(u,v)\leq 2t$, then $f(v)\in B_{L}^{f}(u,2t)$. We have $\tau_{f}^{L}(u)\neq\tau_{f}^{L}(v)$. Thus $C_{\tau_{f}^{L}(u)}\neq C_{\tau_{f}^{L}(v)}$. We have,
@@
Conversely, let $c=(c_{1},c_{2},\dots,c_{s-1})$ and $d=(d_{1},d_{2},\dots,d_{s-1})$ be two codewords of length $s-1$. Then, as the maximum Lee weight of an element in $\mathbb{Z}_{m}$ can be $\left\lfloor\frac{m}{2}\right\rfloor$, we have
@@ 6 FCLMCs for the Lee weight and Lee weight distribution functions
[…]
@@
Observe that $-\left\lfloor\frac{m}{2}\right\rfloor\leq i-j\leq\left\lfloor\frac{m}{2}\right\rfloor$, thus by Eq.2.1, $w_{L}(i-j)=|i-j|$. Hence $d_{L}(u_{i},u_{j})=|i-j|.$
@@
[…]
@@
Note that $(a-1)\left\lfloor\frac{m}{2}\right\rfloor)<i\leq a\left\lfloor\frac{m}{2}\right\rfloor$. This implies that $-\left\lfloor\frac{m}{2}\right\rfloor\leq i-a\left\lfloor\frac{m}{2}\right\rfloor\leq 0\leq\left\lfloor\frac{m}{2}\right\rfloor$. Therefore, again by Eq. 2.1,
@@
$[D_{w_{L}}(t)]_{ij}=\begin{cases}0&\text{ if }i=j\\ [2t+1-|i-j|]^{+}&\text{ if }i\neq j.\end{cases}$

### v2

@@ Abstract
Function-correcting codes are a coding framework designed to minimize redundancy while ensuring that specific functions or computations of encoded data can be reliably recovered, even in the presence of errors. The choice of metric is crucial in designing such codes, as it determines which computations must be protected and how errors are measured and corrected. Previous work by Liu and Liu [6] studied function-correcting codes over $\mathbb{Z}_{2^{l}},\ l\geq 2$ using the homogeneous metric, which coincides with the Lee metric over $\mathbb{Z}_{4}$. In this paper, we extend the study to codes over $\mathbb{Z}_{m},$ for any positive integer $m\geq 2$ under the Lee metric and aim to determine their optimal redundancy. To achieve this, we introduce irregular Lee distance codes and derive upper and lower bounds on the optimal redundancy by characterizing the shortest possible length of such codes. These general bounds are then simplified and applied to specific classes of functions, including Lee-local functions, Lee weight functions, and Lee weight distribution functions. We extend the bounds established by Liu and Liu [6] for codes over $\mathbb{Z}_{4}$ in the Lee metric to the more general setting of $\mathbb{Z}_{m}$. Additionally, we explicitly derive a Plotkin-like bound for linear function-correcting codes in the Lee metric. As the Lee metric coincides with the Hamming metric over the binary field, we demonstrate that our bound naturally reduces to a Plotkin-type bound for function-correcting codes under the Hamming metric over $\mathbb{Z}_{2}$. Furthermore, when the underlying function is bijective, function-correcting codes reduce to classical error-correcting codes. In parallel, our bound correspondingly reduces to the classical Plotkin bound for error-correcting codes, both for the Lee metric over $\mathbb{Z}_{m}$ and for the Hamming metric over $\mathbb{Z}_{2}$.
Keywords: Function-correcting codes, error-correcting codes, Lee metric, optimal redundancy, locally bounded function.
@@ 1 Introduction
Lenz et al. [5] first introduced the framework of function-correcting codes (FCCs) in the Hamming metric, establishing the foundational theory for correcting function evaluations rather than raw symbols. This framework was later extended to other metrics to capture different error models. Specifically, Xia et al. [16] generalized FCC to the symbol-pair metric, while Singh et al. [13] extended it further to the $b$-symbol metric. Building on these developments, in [14], the authors introduced locally $b$-symbol codes and investigated the redundancy of the corresponding function-correcting $b$-symbol codes. Several related advances complement these efforts: Ge et al. [2] studied optimal redundancy bounds for such codes, Ly et al. [7] proposed a conjecture concerning their asymptotic performance, and Premlal et al. [8] developed a function correcting framework for linear functions. In the context of locality, Rajput et al. [9] investigated locally function-correcting codes for the Hamming metric, while Sampath et al. [11] contributed further generalizations of FCC to specialized coding scenarios. Together, these works form a cohesive progression from classical FCC in the Hamming setting to rich variants supporting diverse metrics, locality, and function-specific error corrections.
@@
Function-correcting codes with respect to the homogeneous metric were proposed in [6], where the authors introduced function-correcting codes with homogeneous distance (FCCHDs) as a generalization of function-correcting codes under the Hamming metric. Since both Hamming weight and Lee weight over $\mathbb{Z}_{4}$ can be regarded as special cases of homogeneous weights, the results in [6] naturally extend to the Lee metric over $\mathbb{Z}_{4}$. Motivated by this fact, we investigate the construction and properties of function-correcting codes for the Lee metric over $\mathbb{Z}_{m}$, where $m\geq 2$ is an arbitrary positive integer. We derive upper and lower bounds on the optimal redundancy of function-correcting Lee metric codes for arbitrary functions by employing shortest-length irregular Lee distance codes in Section 3. In Section 4, we provide several bounds on the shortest-length irregular Lee distance codes analogous to the Plotkin and Gilbert-Vashamov bounds. Furthermore, we introduce the concept of locally bounded Lee functions, and using general theoretical results, establish bounds on the optimal redundancy for this class of functions in Section 5. In particular, we demonstrate that the lower bound on optimal redundancy is achievable for locally binary Lee functions. Additionally, we investigate the locality properties of specific functions, the Lee weight function and the Lee weight distribution function, and determine bounds on their optimal redundancy in Section 6. The bounds on the optimal redundancy of function-correcting codes can be obtained as a direct consequence of our general results. In particular, by substituting $m=2$ and $m=4$ into Lemma 4.1, we get the specific cases corresponding to codes in the Hamming distance over $\mathbb{Z}_{2}$ in [5] and the homogeneous distance over $\mathbb{Z}_{4}$ in [6], respectively. This shows that the well-known bounds for these two settings follow naturally from our general framework, without requiring any additional arguments. In Section 7, we investigate function-correcting codes associated with linear functions and derive a Plotkin-type bound on the optimal redundancy. We further demonstrate that this bound generalizes several existing Plotkin-like bounds, each of which emerges as a special case of our result.
@@ 2 Preliminaries
In this section, we discuss basic definitions and results that will be used throughout the paper. We denote the residue ring modulo $m$ by $\mathbb{Z}_{m}$, where $m$ is a positive integer greater than or equal to $2$. We consider the elements of $\mathbb{Z}_{m}$ as $0,1,\dots,m-1$ unless otherwise specified. For a positive integer $k$, $\mathbb{Z}_{m}^{k}$ denotes the set of all $k$-tuples over $\mathbb{Z}_{m}$.
@@
[…]
@@
Lemma 2.3.
[15] Let $\alpha\in\mathbb{Z}_{m}$. Then
[…]
Lemma 2.4.
[15] Let $C=\{C_{1},C_{2},\dots,C_{M}\}\subseteq\mathbb{Z}_{m}^{r}$ be a code of length $r$ and size $M$. Then
[…]
@@ 3 Function-correcting Lee metric codes
Any classical error-correcting code with the minimum Lee distance $2t+1$, defines an $(f,t)$-FCLMC for any function $f:\mathbb{Z}_{m}^{k}\to Im(f)$. For an integer $n$, we denote $n]=\{1,2,\dots,n\}$ and $n]^{+}=\max\{n,0\}$. Define the distance matrix of a given function $f$ as follows.
@@
$D_{f}^{L}(t,u_{1},u_{2},\dots,u_{M})]_{ij}=\begin{cases}[2t+1-d_{L}(u_{i},u_{j})]^{+}&\text{if }f(u_{i})\neq f(u_{j})\\ 0&\text{ otherwise}\end{cases}$
@@
Let $N_{L}(D)$ be the smallest integer $r$ such that there exists a $D$-code of length $r$. When $D]_{ij}=d$ for all $i\neq j$, we denote $N_{L}(M,d)$. Now, we prove a fundamental relationship between the redundancy of optimal FCLMCs and the irregular distance codes.
@@
$D_{f}^{L}(t,e_{1},e_{2},\dots,e_{\eta})]_{ij}=\begin{cases}[2t+1-d_{L}(e_{i},e_{j})]^{+}&\text{ if }i\neq j\\ 0&\text{otherwise}.\end{cases}$
@@ 4 Bounds on the shortest length of irregular-distance codes
[…]
@@
Let $N_{L}(D)=r$. Let $C\subseteq\mathbb{Z}_{m}^{r}$ be a $D$-code of length $r$ and size $M$. Consider an $M\times r$ matrix $A$ whose rows are codewords $\{C_{1},C_{2},\dots,C_{M}\}=C$. By [15, Eq. (15)], we have
$\begin{split}\sum_{1\leq i<j\leq M}d_{L}(C_{i},C_{j})&\leq\frac{SM^{2}r}{2m},\end{split}$
where
[…]
@@
$r\geq\frac{2m}{SM^{2}}\sum_{i,j;i<j}[D]_{ij}.$
This completes the proof. ∎
It is well known that the Lee metric coincides with the Hamming metric over $\mathbb{Z}_{2}$ and coincides with the homogeneous metric over $\mathbb{Z}_{4}$. Hence, we state the lemma for these two metrics as derived in [5] and [6], respectively.
@@
[6] Let $m=4$. For any Lee distance matrix $D\in\mathbb{N}_{0}^{M\times M}$, we have
$N_{L}(D)\geq\frac{2}{M^{2}}\sum_{i,j;i<j}[D]_{ij}.$
@@
[5] Let $m=2$. For any Lee distance matrix $D\in\mathbb{N}_{0}^{M\times M}$, we have
$N_{L}(D)\geq\frac{4}{M^{2}}\sum_{i,j;i<j}[D]_{ij}.$
@@
Let $D$ be a distance matrix with $D]_{ij}=d$ for all $i\neq j$. Then $N_{L}(M,d)\leq N_{H}(M,d)$.
@@ 5 FCLMCs for the locally functions
Let $f:\mathbb{F}_{2}^{k}\to Im(f)$ be a locally $(\lambda,\rho)_{L}$-function. Then there exists a map $\tau_{f}^{L}:\mathbb{F}_{2}^{k}\to[\lambda]$ such that for all $u,v\in\mathbb{F}_{2}^{k}$, if $f(u)\neq f(v)$ and $d_{L}(u,v)\leq\rho$, then $\tau_{f}^{L}(u)\neq\tau_{f}^{L}(v)$
@@
By Lemma 5.3, there exists a map $\tau_{f}^{L}:\mathbb{Z}_{m}^{k}\to[\lambda]$ such that for $u,v\in\mathbb{F}_{2}^{k}$, $\tau_{f}^{L}(u)\neq\tau_{f}^{L}(v)$ whenever $f(u)\neq f(v)$ with $d_{L}(u,v)\leq 2t$. Let $C=\{C_{1},C_{2},\dots,C_{\lambda}\}$ be a code of length $N_{L}(\lambda,2t)$ with $\lambda$ codewords and minimum Lee distance $2t$. Define an encoding $\mathcal{E}:\mathbb{Z}_{m}^{k}\to\mathbb{Z}_{m}^{k+N_{L}(\lambda,2t)}$ given by $\mathcal{E}(u)=(u,C_{\tau_{f}^{L}(u)})$. We claim that this encoding defines an $(f,t)$-FCLMC. Let $u,v\in\mathbb{Z}_{m}^{k}$ with $f(u)\neq f(v)$. Then
@@
Case 2: If $d_{L}(u,v)\leq 2t$, then $\tau_{f}^{L}(u)\neq\tau_{f}^{L}(v)$. Thus $C_{\tau_{f}^{L}(u)}\neq C_{\tau_{f}^{L}(v)}$. We have,
@@
Conversely, let $c=(c_{1},c_{2},\dots,c_{s-1})$ and $d=(d_{1},d_{2},\dots,d_{s-1})$ be two codewords of length s1s-1. Then, as the maximum Lee weight of an element in $\mathbb{Z}_{m}$ can be $\left\lfloor\frac{m}{2}\right\rfloor$, we have
@@ 6 FCLMCs for the Lee weight and Lee weight distribution functions
$d_{L}(u_{i},u_{j})=$ $w_{L}\left(0,0,\dots,i-(a-1)\left\lfloor\frac{m}{2}\right\rfloor-(j-(a-1)\left\lfloor\frac{m}{2}\right\rfloor(a^{\mathrm{th}}\text{-place}),0,\dots,0\right)$
$=$ $w_{L}(0,0,\dots,i-j(a^{\mathrm{th}}\text{-place}),0,\dots,0)$
@@
Observe that $\left\lfloor\frac{m}{2}\right\rfloor\leq i-j\leq\left\lfloor\frac{m}{2}\right\rfloor$, thus by Eq.2.1, $w_{L}(i-j)=|i-j|$. Hence $d_{L}(u_{i},u_{j})=|i-j|.$
@@
$\left(0,0,\dots,0,i-(a-1)\left\lfloor\frac{m}{2}\right\rfloor-\left\lfloor\frac{m}{2}\right\rfloor(a^{\mathrm{th}}\text{-place}),-\left\lfloor\frac{m}{2}\right\rfloor,\dots,-\left\lfloor\frac{m}{2}\right\rfloor,-j+(b-1)\left\lfloor\frac{m}{2}\right\rfloor(b^{\mathrm{th}}\text{-place}),0,\dots,0\right).$
@@
Note that $(a-1)\left\lfloor\frac{m}{2}\right\rfloor)<i\leq a\left\lfloor\frac{m}{2}\right\rfloor$. This implies that $\left\lfloor\frac{m}{2}\right\rfloor\leq i-a\left\lfloor\frac{m}{2}\right\rfloor\leq 0\leq\left\lfloor\frac{m}{2}\right\rfloor$. Therefore, again by Eq. 2.1,
@@
$D_{w_{L}}(t)]_{ij}=\begin{cases}0&\text{ if }i=j\\ [2t+1-|i-j|]^{+}&\text{ if }i\neq j.\end{cases}$
@@
7 Plotkin-like bound for linear function-correcting codes in the Lee metric
In this section, we derive a Plotkin-like bound for a specific class of functions, namely linear functions, in the context of channels measured using the Lee metric. Linear function correcting codes over the Hamming metric are discussed in [8], and later extended to $b$-symbol read channels in [11].
Theorem 7.1.
Let $f:\mathbb{Z}_{m}^{k}\to\mathbb{Z}_{m}^{\ell}$ be an onto linear function. Then the optimal redundancy of an $(f,t)$-FCLMC is bounded by
$r_{f}^{L}(k,t)\geq\frac{m}{S}(2t+1)(1-m^{-\ell})-k+\frac{s}{Sm^{k-1}},$
where $S$ is as in Lemma 2.3 and $s=\sum_{u\in\ker(f)}w_{L}(u)$.
Proof.
Let $M=m^{k}$, $\mathbb{Z}_{m}^{k}=\{u_{1},u_{2},\dots,u_{M}\}$, and $D=D_{f}^{L}(t,u_{1},u_{2},\dots,u_{M})$. Let $C=\{C_{1},C_{2},\dots,C_{M}\}\subseteq\mathbb{Z}_{m}^{r}$ be a $D$-code of length $r$, that is, $d_{L}(C_{i},C_{j})\geq[D]_{ij}$. Then, by Lemma 2.4, we have
$\frac{SM^{2}r}{m}\geq\sum_{i,j}d_{L}(C_{i},C_{j})\geq\sum_{i,j}[D]_{ij}.$ (7.1)
Since $f$ is an onto linear function, therefore cardinality of $\ker(f)$ is $m^{k-\ell}$. Thus, each coset of $\ker(f)$ in $\mathbb{Z}_{m}^{k}$ contains $m^{k-\ell}$ elements. Observe that if $u_{i},u_{j}$ belong to the same coset, then $f(u_{i})=f(u_{j})$. Consequently, $D]_{ij}=0$, that is, each column of $D$ contains at least $m^{k-\ell}$ zeros. When $u_{i},u_{j}$ are in different cosets, then $D]_{ij}=[2t+1-d_{L}(u_{i},u_{j})]^{+}$. Let $0\neq v\in\mathbb{Z}_{m}^{k}$, $u_{i^{\prime}}=u_{i}+v$ and $u_{j^{\prime}}=u_{j}+v$. Then $d_{L}(u_{i},u_{j})=d_{L}(u_{i^{\prime}},u_{j^{\prime}})$. Therefore, each column (row) in $D$ contains the same set of entries, implying that the columns are permutations of one another. Thus,
$\sum_{i,j}[D]_{ij}=(\text{no. of columns)}\times(\text{sum of one column)}.$
Let $u_{1}=\bm{0}$ be the zero vector in $\mathbb{Z}_{m}^{k}$. We calculate the sum of the entries of the column corresponding to $u_{1}$. Let $I=\mathbb{Z}_{m}^{k}\setminus\ker(f)$. If $u_{i}\in\ker(f)$, then $D]_{i1}=0$. For $i\in I$, $D]_{i1}\geq 2t+1-d_{L}(u_{i},\bm{0})=2t+1-w_{L}(u_{i})$. Thus,
$\sum_{i}[D]_{i1}\geq(2t+1)(m^{k}-m^{k-\ell})-\sum_{i\in I}w_{L}(u_{i}).$
The sum of Lee weights of all the vectors in $\mathbb{Z}_{m}^{k}$ is $kSm^{k-1}$. Thus,
$\sum_{i\in I}w_{L}(u_{i})=kSm^{k-1}-s,\text{ where }s=\sum_{u\in\ker(f)}w_{L}(u).$
Therefore,
$\begin{split}\sum_{i,j}[D]_{ij}&\geq M\cdot\sum_{i}[D]_{i1}\\ &\geq m^{k}((2t+1)(m^{k}-m^{k-\ell})-kSm^{k-1}+s.\end{split}$ (7.2)
The desired result follows from Eqs. 7.1 and 7.2. ∎
As stated in Lemma 4.1, the Plotkin bound depends on the sum of all $D]_{ij}$ terms. While in the case of linear functions, this reduces to the sum of the Lee weights of the vectors in $\ker(f)$. In [8], the authors derived a Plotkin-like bound for linear function-correcting codes over the finite field $\mathbb{F}_{q}$ under the Hamming metric. As the Lee metric coincides with the Hamming metric over $\mathbb{Z}_{2}$, our corresponding result in the Hamming metric over $\mathbb{Z}_{2}$ is consistent with the bound in [8].
Corollary 7.2.
Let $f:\mathbb{Z}_{2}^{k}\to\mathbb{Z}_{2}^{\ell}$ be an onto linear function. Then the optimal redundancy of an $(f,t)$-FCC is bounded by
$r_{f}^{H}(k,t)\geq 2(2t+1)(1-2^{-\ell})-k+\frac{s}{2^{k-1}},$
where $s=\sum_{u\in\ker(f)}w_{H}(u)$.
Proof.
By Lemma 2.3, we have $S=1$ for $m=2$. The rest follows from Theorem 7.1 by putting $m=2$. ∎
When the function $f$ is bijective, a function-correcting code reduces to a systematic error-correcting code. Particularly, setting $f$ to be a bijective function, that is, $s=0$ and $\ell=k$, the bound presented in Theorem 7.1 simplifies to the Poltkin bound for error-correcting codes in the Lee metric. This aligns precisely with the well-known Plotkin low-rate average distance bound (see [15]).
Corollary 7.3.
Let $f:\mathbb{Z}_{m}^{k}\to\mathbb{Z}_{m}^{k}$ be a bijective linear function. Then the optimal redundancy of an $(f,t)$-FCLMC is bounded by
$n=r_{f}^{L}(k,t)+k\geq\frac{m}{S}(2t+1)\left(\frac{m^{k}-1}{m^{k}}\right),$
where $S$ is as in Lemma 2.3.
The following corollary derives the classical Plotkin bound for error-correcting codes over $\mathbb{Z}_{2}$ in the Hamming metric.
Corollary 7.4.
Let $f:\mathbb{Z}_{2}^{k}\to\mathbb{Z}_{2}^{k}$ be a bijective linear function. Then the optimal redundancy of an $(f,t)$-FCC is bounded by
$n=r_{f}^{H}(k,t)+k\geq(2t+1)\left(\frac{2^{k}-1}{2^{k-1}}\right).$